### v1

# Self-supervised Pre-training for Mapping of Archaeological Stone Wall in Historic Landscapes Using High-Resolution DEM Derivatives

Abstract
Dry-stone walls hold significant heritage and environmental value. Mapping and maintaining these structures is essential for preserving ecosystems and supporting wildfire control during dry seasons in Australia. However, a substantial portion of stone walls located in wild landscapes remain unidentified and unregistered due to their inaccessibility and the high cost of mapping. Deep learning-based segmentation models offer a scalable solution for automated mapping of dry-stone walls, but two challenges remain, including (1) visual occlusion of low-lying stone walls by dense vegetation and (2) limited availability of labeled data for training supervised segmentation models. In this study, we propose a segmentation approach for the automatic and accurate mapping of low-lying dry-stone walls using high-resolution digital elevation models (DEMs). The use of Airborne LiDAR-derived DEMs help overcome the visual occlusion problem caused by dense vegetation and canopy cover, by capturing underlying terrain structures that are invisible in traditional optical imagery. Our approach proposes DINO-CV, a self-supervised cross-view pre-training strategy based on knowledge distillation, to address the challenge of limited labeled data. DINO-CV learns invariant visual and structural features across multiple high-resolution DEM derivatives, enabling accurate stone wall mapping on various computer vision backbones (i.e., ResNet, Wide ResNet, and Vision Transformers (ViTs)). Applied to the UNESCO World Heritage cultural landscape of Budj Bim, Victoria, our method successfully identifies one of Australia’s densest collections of colonial dry-stone walls beyond Indigenous cultural heritage. The proposed method achieves a mean Intersection over Union (mIoU) of 68.6% between predicted and ground-truth stone wall locations in the test areas, and maintains robust performance of 63.8% mIoU when fine-tuned with limited labeled data (10%) using self-supervised vision backbones pre-trained on the DEM derivatives of Budj Bim landscape. These results demonstrate the potential of applying self-supervised learning on high-resolution DEM derivatives for automated mapping of dry-stone walls, particularly in densely vegetated environments and heritage landscapes with limited labeled data.
keywords:
Mapping of historic dry-stone walls , Self-supervised learning , DEMs segmentation , Wildfire control
journal: Elsevier
[…]
[inst1]organization=The University of Melbourne,city=Parkville, postcode=3010, state=Victoria, country=Australia
1 Introduction
Dry-stone walls are significant cultural heritage structures that have been constructed for thousands of years in various regions around the world (UNESCO, 2018). These walls are built without mortar, using carefully selected and placed stones to create durable and functional barriers (UNESCO, 2018; Doneus et al., 2022). In Australia, dry-stone walls are often associated with European colonial history and are typically used for agricultural purposes, such as livestock containment and land management (Marshall, 2018). They also hold cultural significance for Indigenous communities, who have their own traditional stone structures and practices (Jiménez de Madariaga, 2021). The mapping and maintenance of dry-stone walls are essential for preserving ecosystems (Preti et al., 2018; Camera et al., 2018) and supporting wildfire control during dry seasons in Australia. However, a substantial portion of stone walls located in wild Indigenous landscapes remain unidentified and undocumented due to limited accessibility and the high cost of conventional mapping approaches.
Traditional methods of mapping dry-stone walls, such as manual surveys or aerial photography, can be impractical and labor-intensive, especially in remote or rugged terrain. Recent advancements in computer vision and deep learning have enabled the development of automated, scalable mapping techniques that can efficiently detect dry-stone walls from high-resolution imagery and LiDAR data (Trotter et al., 2022; Suh and Ouimet, 2023; Wang et al., 2024). Supervised deep neural networks, such as ConvNets (CNNs) (He et al., 2016), vision transformers (ViTs) (Dosovitskiy et al., 2021), and U-Net architectures (Ronneberger et al., 2015), can be trained on annotated datasets to learn the visual and spatial features of dry-stone walls, including their shape, texture, and spatial patterns. These deep-learning based mapping approaches largely reduce the time and cost of mapping and offer the potential for large-scale registration of undocumented structures. However, the performance of these deep supervised models heavily depends on the availability of large, labeled datasets for model training, which are often difficult and costly to obtain. Another critical challenge is the visual occlusion of low-lying stone wall structures by dense vegetation in wild landscapes, which makes it difficult for models to accurately detect and localize these features.
Self-supervised learning (SSL) has emerged as a promising approach to mitigate the challenge of limited labeled data for deep supervised models (Caron et al., 2020, 2021; Chen et al., 2020a; Grill et al., 2020; He et al., 2020). SSL methods are capable of learning object-centric visual representations from large unlabeled datasets, enabling the transfer of knowledge to downstream tasks such as image segmentation and classification. By pre-training models on large unlabeled datasets, SSL can improve the performance of supervised models trained with limited labeled data. In the context of dry-stone wall mapping, SSL is particularly beneficial for learning semantic and geometric features of stone walls from high-resolution aerial imagery or digital elevation models (DEMs) without extensive manual annotation. This is especially relevant in heritage landscapes, where the availability of labeled data may be limited due to the complexity and variability of the structures.
To overcome the challenge of visual occlusion of low-lying stone walls by dense vegetation, the use of high-resolution DEMs and their derivatives (i.e., multi-directional hillshadings (MHS) and Visualization for Archaeological Topography (VAT)) for stone wall mapping can provide rich topographic information that enhances the visibility of terrain features in complex, vegetated environments. These DEM derivatives are typically processed from airborne LiDAR data, which can penetrate dense vegetation canopies to capture detailed ground surface information. In contrast, image-based data often struggle to generate accurate DEMs in vegetated locations where the terrain is heavily obscured (White et al., 2013).
Grounded in these hypotheses, this study poses the following research question: How can self-supervised learning on high-resolution DEM derivatives improve the mapping performance of low-lying dry-stone walls in complex, vegetated environments? To answer this question, we explore the potential of self-supervised learning on high-resolution DEM derivatives for automated mapping of dry-stone walls in complex, vegetated environments. Hence, we propose a self-supervised cross-view pre-training strategy, DINO-CV, based on knowledge distillation to learn visual and structural features from multiple DEM derivatives generated from high-resolution aerial LiDAR data. The mapping performance of our approach is evaluated on the Budj Bim cultural landscape in Victoria, Australia, a UNESCO World Heritage site known for its extensive colonial dry-stone walls and significant Indigenous heritage. We discover that these self-supervised features enable accurate mapping of stone wall structures on various computer vision backbones, including ResNet, Wide ResNet, and Vision Transformers (ViTs), through limited supervised fine-tuning. The proposed self-supervised pre-training strategy, DINO-CV, can learn structural and visual features from different terrain data views (i.e., MHS and VAT), thereby substantially reducing the need for labeled data in supervised fine-tuning. These findings highlight the potential of self-supervised learning on DEM derivatives for large-scale, automated mapping of stone walls in complex, vegetated, and data-scarce heritage environments.
2 Background
2.1 DEM Visualization of Archaeological Landscapes
Detecting archaeological features using high-resolution topographic data has become a central focus in archaeological remote sensing (Chase et al., 2017; Štular et al., 2021). Airborne LiDAR systems provide highly accurate 3D measurements of the landscape by emitting laser pulses that penetrate vegetation, capturing the underlying bare-earth surface. This capability is especially advantageous in densely vegetated regions, where traditional aerial imagery may obscure subtle topographic variations.
These LiDAR-derived measurements can be processed into digital elevation models (DEMs), which represent the elevation of the bare-earth surface by filtering out vegetation and non-ground elements. DEMs can provide detailed topographic information that enables the detection of subtle landscape features, including ancient structures, earthworks, and other archaeological remains. As a result, DEMs have shown particularly suitable for mapping and analyzing large, complex, and vegetated archaeological landscapes (Štular et al., 2012). To further enhance the visibility of these features, a variety of DEM visualization techniques have been developed, such as hillshading (Yoeli, 1967), slope analysis (Doneus and Briese, 2006), and sky-view factor (Kokalj et al., 2011). These methods aid researchers in identifying and interpreting landscape features that are often not discernible in standard aerial or satellite imagery.
One of the DEM visualization methods, Visualization for Archaeological Topography (VAT) developed by Verbovšek et al. (2019); Kokalj and Somrak (2019), integrates multiple topographic layers, including hillshaded relief, slope, positive openness, and sky-view factor, into a single composite image. This combination method enhances the detection of subtle topographic variations that may indicate buried or obscured archaeological features. Another widely used technique, multi-directional hillshading (MHS) (Devereux et al., 2008), generates multiple hillshade images from different illumination angles and combines them into a multi-band raster or composite image. This method improves the visibility of terrain features such as slopes, ridges, and embankments, while reducing shadow saturation and mitigating directional bias in DEMs. For instance, Comer et al. (2019) employed grayscale MHS visualizations derived from LiDAR data to reveal subtle cultivation features, including irrigation berms and channels, on Temwen Island and the islets of Nan Madol.
Together, these visualization techniques serve as powerful tools for archaeological mapping and interpretation, enabling researchers to identify, map, and analyze archaeological features embedded in complex cultural landscapes that might otherwise remain undetected (Guyot et al., 2021b, a). A recent study by Jaturapitpornchai et al. (2024) evaluated the impact of various DEM visualizations on the segmentation performance of archaeological features using deep learning models. The study concludes that identifying a universally optimal visualization method remains inconclusive, as performance varies across archaeological classes and landscape contexts.
2.2 Automated Mapping of Dry-Stone Wall Structures
Recent advances in deep learning have enabled the automated detection and mapping of archaeological features from remote sensing data. Previous study by Trotter et al. (2022) employed a supervised Fully convolutional Neural Networ (FCN) (Long et al., 2015) based on the U-Net architecture (Ronneberger et al., 2015) to automate the mapping of dry-stone walls in Denmark. Their model was trained on a dataset of digital terrain models (DTMs) with a spatial resolution of 0.4m, using ground-truth annotations of dry-stone wall locations. While the study demonstrated the potential of deep learning for this task, its applicability is limited by the availability of high-resolution terrain data, which is widely accessible in Denmark but not available in other regions. Furthermore, the Danish DTM dataset is updated only every five years, and the performance of the supervised model is constrained by the limited quantity of labeled training data.
Another study by Suh and Ouimet (2023) proposed an automated mapping framework based on U-Net (Ronneberger et al., 2015) and ResU-Net (Diakogiannis et al., 2020) architectures for detecting dry-stone walls in forested terrain in the Northeastern United States. Their work investigated the impact of the quality of high-resolution DEM derivatives, specifically hillshade and slope maps, generated from airborne LiDAR data on model performance. They found that illumination parameters, such as sunlight angle and the brightness or darkness of hillshade images, were important factors that affect the model’s performance on detecting small-scale geomorphic features. These findings highlight the importance of optimizing DEM visualization inputs when applying deep learning models to archaeological mapping tasks.
A related study by Wang et al. (2024) developed a segmentation framework for mapping the ancient city walls from airborne LiDAR data in archaeological landscapes. Focusing on the ancient capital of Jinancheng, their method involved converting airborne LiDAR point clouds into DEMs, followed by the application of a U-Net architecture to segment wall remains at the pixel level. The study ensures the quality of labeled data of city wall by expert interpretation based on the archaeological survey data, and increase the number of training samples by data augmentation for model training. A post-processing step using connected component analysis was applied to refine the segmentation outputs of model by removing noise and filling small gaps.
These studies highlight the potential of deep learning for the automated mapping of archaeological stone wall structures. However, a common limitation across existing approaches is the dependence on large quantities of high-quality labeled data to achieve strong segmentation performance. The generation of such datasets often requires expert knowledge and significant manual effort, which can be resource-intensive. Moreover, the limited availability of high-resolution terrain data in many regions restricts the generalizability and scalability of these methods to diverse archaeological contexts.
2.3 Self-supervision with Knowledge Distillation
Self-supervised learning (SSL) methods based on knowledge distillation have been proposed as an effective approach for learning object-centric visual features from large, unlabeled datasets, enabling improved performance on downstream tasks such as image segmentation and classification (Grill et al., 2020; Caron et al., 2021). By pre-training models on large, unlabeled datasets, SSL can improve the performance of supervised models even with limited labeled data.
An initial study proposed by Hinton et al. (2015), as a method to transfer knowledge from a large ensemble of models (teachers) into a single, compact student model. This approach enables efficient deployment while retaining the performance benefits of model ensembles. Recent SSL methods based on knowledge distillation, such as DINO (Caron et al., 2021) and BYOL (Grill et al., 2020), have adapted this teacher–student framework to a self-supervised, label-free setting by using momentum encoders and self-distillation to match intermediate feature representations and output distributions between teacher and student networks. These methods do not require labeled data and are well-suited for learning semantic and structural features from visual inputs. Recent extensions, DINOv2 (Oquab et al., 2024) and DINOv3 (Siméoni et al., 2025), push this paradigm to much larger scales. DINOv2 stabilizes training at scale and introduces an automated pipeline for curating large and diverse image datasets. DINOv3 further scales the approach to 7B parameters and trains on even larger curated datasets, emphasizing scalability and generalization. While recent extensions such as DINOv2 and DINOv3 push self-distillation toward foundation-scale models through large curated datasets and massive ViTs, these directions are beyond the scope of this work. Instead, our focus is on cross-view self-supervision tailored to high-resolution DEM derivatives, addressing archaeological mapping challenges where labeled data is scarce.
In parallel of SSL with self-distillation, contrastive self-supervised learning methods such as SimCLR (Chen et al., 2020a, b), MoCo (He et al., 2020), and MoCov2 (Chen et al., 2020c) learn visual features by maximizing agreement between different augmented views of the same image (positive pairs), while minimizing agreement between views of different images (negative pairs). This is typically achieved by contrasting representations in a latent embedding space. Although contrastive learning has shown strong performance across a range of visual tasks, it often requires large batch sizes or memory banks to provide a sufficient number of negative examples for effective training (Caron et al., 2021).
Both contrastive and distillation-based SSL frameworks have demonstrated state-of-the-art performance on various vision tasks and are particularly promising for applications in remote sensing and archaeological feature mapping, where labeled data is limited but large volumes of unlabeled imagery or elevation data are available. In the domain of remote sensing, the SeCo framework proposed by Manas et al. (2021) leverages seasonal variations in satellite imagery, along with temporal and spatial invariance constraints to learn robust and temporal-invariant visual representations through contrastive learning. The knowledge distillation framework, DINO-MM (Wang et al., 2022), extends the DINO framework to multimodal data, enabling the learning of joint representations from different modalities (i.e., optical and SAR imagery). This approach has demonstrated strong performance in remote sensing tasks such as image classification and retrieval. DINO-MC (Wanyan et al., 2024) introduces a multi-size cropping strategy to strengthen global–local view alignment between teacher and student networks. By incorporating local crops of varied sizes, the model learns more diverse and scale-invariant visual features on the SeCo-100K (Manas et al., 2021) dataset, which enhances its performance across a range of downstream tasks, including remote sensing imagery classification and change detection.
Differently, we propose a self-supervised cross-view pre-training framework based on knowledge distillation paradigm to pre-train vision backbones on unlabeled DEM derivatives (i.e., MHS and VAT), which can improve stone wall mapping capability in complex and vegetated archaeological landscapes.
3 Methodology
Figure 1: The aerial image (Row 1), Multi-directional Hillshade (MHS, Row 2), and Visualization for Archaeological Topography (VAT, Row 3) of dry-stone walls in the Budj Bim landscape. From left to right, columns show: (1–2) multiple parallel thin walls and an isolated thin wall; (3–4) thick, zigzag walls and a straight wall partially occluded by vegetation; (5–6) roadside straight walls with gate openings, and a heavily occluded wall segment.
3.1 DEM Visualization of Budj Bim Landscapes
High-resolution DEM visualization techniques are useful visualization tools for mapping cultural heritage features in complex terrains, where archaeological structures such as dry-stone walls are often obscured in conventional aerial imagery. To enhance the visibility of low-lying structures and support automated stone wall mapping on archaeological landscape, we employ two DEM-derived visualization techniques: Multi-directional Hillshade (MHS) and Visualization for Archaeological Topography (VAT).
MHS simulates illumination from multiple sun angles to enhance the perception of topographic features, such as ridges, edges, and linear structures embedded in the landscape. This visualization technique provides a general-purpose view of terrain geometry, improving the contrast and continuity of linear archaeological features. In contrast, VAT emphasizes micro-topographic variations by combining multiple elevation derivatives, including slope, local relief, openness, and sky-view factor, into a single composite image. VAT is designed specifically to highlight subtle archaeological features such as buried walls, ditches, and mounds that may not be visible in traditional hillshading.
While aerial imagery provides high-fidelity semantic information, such as textures of objects and surrounding vegetation, it is often limited by occlusions caused by dense vegetation, particularly in rural or forested areas. In contrast, MHS and VAT visualizations derived from DEM data are unaffected by such occlusions and can effectively reveal dry-stone wall structures that are otherwise indiscernible in aerial views, as illustrated in Fig. 1. In detail, columns 1 and 2 of Fig. 1 highlight multiple thin, parallel stone walls and a solitary linear segment, which are clearly discernible in both MHS and VAT. Columns 3 and 4 display zigzag and straight thick walls that are partially or fully concealed by vegetation in the aerial view but remain visible in the DEM visualizations. Columns 5 and 6 show roadside walls with gate openings and heavily occluded segments.
Together, MHS and VAT offer complementary topographic features that are well-suited for detecting the diverse forms and different kinds of dry-stone walls across the Budj Bim cultural landscape. Their integration enables robust feature extraction and significantly enhances automated mapping performance where visual occlusion and subtle relief variations pose challenges to traditional image-based methods.
3.2 Self-supervised Cross-view Pre-training
Figure 2: (a) The DINO-CV framework for self-supervised pre-training. At each training iteration $k$, the student network $g_{\theta_{s}}$ receives a randomly sampled input from the collection of DEM-derived views (i.e., MHS and VAT), while the teacher network $g_{\theta_{t}}$ receives a fixed view (MHS). (b) The Pseudo Siamese Network used for supervised fine-tuning. This architecture comprises two teacher networks, each pre-trained with DINO-CV on a distinct input view (MHS and VAT), enabling the extraction of complementary features for stone wall mapping.
Knowledge Distillation
The self-supervised cross-view pre-training framework, DINO-CV (Fig.2(a)), is designed to learn invariant visual and structural features across multiple high-resolution DEM derivatives to enhance mapping performance of models in the subsequent supervised fine-tuning. DINO-CV builds on the knowledge distillation paradigm introduced by Caron et al. (2021), where a student network $g_{\theta_{s}}$, comprising a vision backbone $f_{\theta_{s}}$ and a projection head $h_{\theta_{s}}$, such that $g_{\theta_{s}}=h_{\theta_{s}}\circ f_{\theta_{s}}$, is trained to match the output distribution of a teacher network $g_{\theta_{t}}$ by minimizing the cross-entropy between their temperature-scaled softmax outputs:
$\min_{\theta_{s}}\;\mathcal{L}\left(P_{t}(g_{\theta_{t}}(x)),P_{s}(g_{\theta_{s}}(x))\right),$ (1)
where $P_{t}$ and $P_{s}$ are the softmax-normalized outputs of the teacher and student networks, respectively.
During training, the student parameters $\theta_{s}$ are updated using stochastic gradient descent (SGD) to minimize the distillation loss. The teacher network however is not directly optimized; instead, its parameters $\theta_{t}$ are updated at each iteration using an exponential moving average (EMA) of the student’s parameters (Tarvainen and Valpola, 2017):
$\theta_{t}\leftarrow m\theta_{t}+(1-m)\theta_{s},$ (2)
where $m\in[0,1)$ is a momentum coefficient that controls the smoothness of the teacher update.
To prevent model collapse, the output of the teacher network is centered by adding a running mean vector $c$ before softmax normalization:
$g_{\theta_{t}}(x)\leftarrow g_{\theta_{t}}(x)+c.$
The center $c$ is updated at each iteration based on the current batch of teacher outputs:
$c\leftarrow mc+(1-m)\frac{1}{B}\sum_{i=1}^{B}g_{\theta_{t}}(x_{i}),$ (3)
where $B$ is the batch size. In addition, output sharpening is applied by using a low temperature $\tau_{t}$ in the teacher softmax, which concentrates the probability mass and encourages the student to produce confident predictions.
Cross-view Pre-training
The core idea of DINO-CV is learning invariant visual and structural terrain features across multiple DEM derivatives (e.g., MHS and VAT). In the cross-view pre-training strategy, the student network is trained to learn view-invariant features by matching stable targets derived from varying data inputs. Let $\mathcal{V}=\{v_{1},...,v_{n}\}$ be a collection of DEM derivatives (e.g., MHS and VAT), and let $x^{(i)}\sim\mathcal{V}$ be a randomly sampled data view for the student, while $x^{(t)}$ denotes a fixed view (e.g., MHS) for the teacher.
Following the multi-crop augmentation strategy of Caron et al. (2021), multiple global and local crops are generated from both inputs: $v_{s}=\text{aug}(x^{(i)})$ and $v_{t}=\text{aug}(x^{(t)})$. These crops are passed into the networks to compute a set of representations from the student: $\{g_{\theta_{s}}(v^{c}_{s})\}_{c=1}^{N}$, where $N$ is the number of student crops (global + local), and a set of global representations from the teacher: $\{g_{\theta_{t}}(v^{c}_{t})\}_{c=1}^{M}$, where $M$ is the number of teacher global crops. The student is trained to match the teacher’s output distributions by minimizing the cross-entropy loss in Eq. 1 across all teacher-student crop pairs.
Convergence of the Teacher Network
In the pre-training of DINO-CV, while the teacher network $g_{\theta_{t}}$ receives only a fixed data view (e.g., MHS or VAT) as input, the EMA update (Eq. 2) ensures that the teacher progressively inherits the learned cross-view invariance and semantically meaningful representations from the student network $g_{\theta_{s}}$, as follows:
$\theta_{t}^{(k+1)}=m\theta_{t}^{(k)}+(1-m)\theta_{s}^{(k+1)},$ (4)
where $m\in[0,1)$ is a momentum coefficient controlling the smoothness of parameters update.
At each iteration $k$, the student parameters $\theta_{s}^{(k)}$ are updated by SGD with a randomly sampled data view $x^{(i_{k})}\sim\mathcal{V}$ from the collection of DEM derivatives:
$\theta_{s}^{(k+1)}=\theta_{s}^{(k)}-\eta\nabla_{\theta_{s}}\mathcal{L}_{k},$ (5)
where $\eta$ is the learning rate for weight update, and $\mathcal{L}_{k}$ is the cross-entropy loss between the teacher’s output on the fixed view $x^{(t)}$ and the student’s output on the random view $x^{(i_{k})}$. This however introduces non-stationarity into the training process, as the student’s input view distribution changes over time.
Despite this randomness in student inputs, the teacher can still converge, assuming that the distillation loss $\mathcal{L}_{k}$ consistently aligns the student’s outputs to stable targets. Since the teacher’s output is anchored to a fixed input data view and the student is repeatedly trained to match this output under varying inputs, the student gradually learns to produce view-invariant representations. Consequently, the teacher accumulates and smooths these representations via the EMA of student weights.
By unrolling the EMA update from Eq. 4, the teacher’s parameters update at iteration $k$ can be expressed as:
$\theta_{t}^{(k)}=(1-m)\sum_{i=1}^{k}m^{k-i}\theta_{s}^{(i)}+m^{k}\theta_{t}^{(0)},$ (6)
which is a weighted moving average of the past student parameters.
Assuming that the student parameters converge to a stable point:
$\lim_{k\to\infty}\theta_{s}^{(k)}=\theta^{*},$
as expected when the loss $\mathcal{L}_{k}$ decreases and the randomly varying input data views are semantically aligned, the teacher parameters also converge:
$\lim_{k\to\infty}\theta_{t}^{(k)}=\theta^{*}.$
This ensures that the teacher network generates stable, smooth, and semantically consistent targets throughout the cross-view training process.
3.3 Pseudo Siamese Network
In supervised fine-tuning, we adopt a Pseudo Siamese Network architecture (Fig. 2(b)) to effectively leverage the complementary representations of DEM derivatives. This architecture consists of two parallel teacher networks pre-trained with DINO-CV, each trained on a different data view. A Pseudo Siamese Network refers to a two-stream architecture where the networks share the same architecture but have independent weights. In our case, the MHS and VAT streams are parameterized by their respective teacher networks, each pre-trained separately with DINO-CV, allowing the two streams to specialize in their respective data view.
Let $g_{\theta_{t}^{\text{MHS}}}$ and $g_{\theta_{t}^{\text{VAT}}}$ denote the two pre-trained teacher networks: the teacher network trained with MHS inputs $g_{\theta_{t}^{\text{MHS}}}(x^{\text{MHS}})$ and with VAT inputs $g_{\theta_{t}^{\text{VAT}}}(x^{\text{VAT}})$.
During supervised fine-tuning, each data-view-specific input is passed into its corresponding teacher network, and their latent representations are fused via element-wise summation:
$z(x)=g_{\theta_{t}^{\text{MHS}}}(x^{\text{MHS}})+g_{\theta_{t}^{\text{VAT}}}(x^{\text{VAT}}).$ (7)
The fused representation $z(x)$ is then passed to a decoder head $Dec(z(x))$ for pixel-wise semantic segmentation of dry-stone walls. Depending on the encoder backbone, we use: 1. U-Net decoder (Ronneberger et al., 2015) for ConNet backbones (e.g., ResNet-50 and Wide ResNet-50), and 2. DPT decoder (Ranftl et al., 2021) for ViT backbones (e.g., ViT-Small/16). This architectural design also benefits from the view-invariant representations learned by the student network during cross-view pre-training. Since the parameters of teacher networks are derived from the student networks trained to align outputs across randomly sampled data views, the fused representations $z(x)$ encode robust, view-invariant structural features transferable across different data views.
3.4 Implementation
Pre-training
We pre-train three widely used vision backbones: ResNet-50 (He et al., 2016), Wide-ResNet-50-2 (Zagoruyko and Komodakis, 2016), and ViT-S/16 (Dosovitskiy et al., 2021), within the DINO-CV framework. The choice of ResNet-50 and ViT-S/16 is motivated by their comparable model sizes (21M vs. 22M parameters) and contrasting architectural paradigms (convolutional vs. attention-based), enabling a fair comparison of ConvNets and Transformers for the task of dry-stone wall mapping. Wide-ResNet-50-2 is chosen for its stronger learning capacity through a wider architecture, and has demonstrated improved performance across various vision tasks.
Pre-training follows the implementation details of DINO (Caron et al., 2021). ViT-S/16 is optimized using AdamW (Loshchilov and Hutter, 2019), while ResNet-50 and Wide-ResNet-50-2 are trained using SGD with momentum. All models are trained with a batch size of 256 for 100 epochs, distributed over two NVIDIA A100 GPUs. The learning rate is linearly warmed up during the first 10 epochs to an initial value of 0.0005 for ViT-S/16 and 0.001 for the ResNet variants, followed by a cosine decay schedule. The base momentum parameter $m$ for the EMA update (Eq. 2) of the teacher network starts at 0.996 and is gradually increased to 1.0 following a cosine schedule.
We apply data augmentations consistent with those used in DINO (Caron et al., 2021) and BYOL (Grill et al., 2020), including the local and global multi-crop strategy. For ViT-S/16, global crops are randomly sampled from scale ranges $(0.33,1.0)$ and local crops from $(0.1,0.33)$. For ResNet-based models, global crops are randomly sampled from $(0.15,1.0)$ and local crops from $(0.05,0.15)$.
Fine-tuning
We fine-tune the pre-trained vision backbones for 100 epochs on a labeled dataset (Sec 4.1) for stone wall mapping. ResNet-based models are optimized using SGD with a batch size of 64, an initial learning rate of 0.01, and a weight decay of 0.0001. For the ViT-S/16 backbone, we employ the AdamW optimizer with a batch size of 64, a learning rate of 0.00005, and a weight decay of 0.05. For data augmentation in training, input data is resized into $256^{2}$ using bicubic interpolation, followed by random rotations up to 180 degrees. Image normalization is performed according to ImageNet statistics (Deng et al., 2009). For mask inputs, resizing and rotations are performed using nearest neighbor interpolation to preserve label integrity.
4 Experiments
4.1 Dataset
Table LABEL:tab:datasets provides a comprehensive summary of the key specifics of the datasets used in this study.
BudjBimLandscape
For self-supervised pre-training, we apply a dataset called BudjBimLandscape, covering 301 km2 of terrain in the Budj Bim cultural landscape, located in southeastern Australia. The dataset comprises unlabeled aerial imagery and two DEM-derived views: multi-directional hillshading (MHS) and visually anisotropic terrain (VAT), both generated from a high-resolution DEM at 50 cm spatial resolution (see Sec. 3.1). Due to the visual occlusion of low-lying stone walls by dense vegetation in aerial imagery, only the MHS and VAT data views are used for DINO-CV pre-training. The dataset contains a total of 188,006 image tiles of size $400^{2}$ pixels, which are resized to $256^{2}$ for training.
BudjBimArea
For supervised fine-tuning and model evaluation, we use an annotated dataset called BudjBimArea, covering a 25 km2 region in the northern part of the Budj Bim landscape. This area features one of the highest densities of European historic dry-stone walls in Australia. The region is spatially divided into six equal rectangular zones, as illustrated in Fig. 3. Wall annotations were manually created using the Semi-Automatic Classification Plugin within QGIS (Congedo, 2021). Annotations are provided as binary raster masks in the $xy$-plane, where a pixel value of 1 denotes the presence of a wall, and 0 denotes background (non-wall) regions. The dataset is publicly released to support reproducibility and benchmarking.
The dataset comprises a total of 7,102 image tiles, each with a resolution of $400^{2}$ pixels, which are resized to $256^{2}$ for training and evaluation. We adopt a six-fold leave-one-area-out cross-validation strategy, where in each fold, five of the six areas are used for training and validation, and the remaining area is held out for testing. Within the training subset, data is further split into 80% for training and 20% for validation. This splitting strategy ensures evaluation across diverse spatial regions of dataset and provides a robust assessment of generalization to unseen regions.
Table 1: Summary of the BudjBimLandscape and BudjBimArea datasets used for self-supervised pre-training and supervised fine-tuning.
Description BudjBimLandscape BudjBimArea
Usage Self-supervised pre-training Supervised fine-tuning and evaluation
Coverage Area 301 km2 25 km2
Location Budj Bim Cultural Landscape, VIC Northern Budj Bim region
Data Views MHS, VAT MHS, VAT, binary wall annotations
DEM Resolution 0.5 m 0.5 m
Tile Size $400\times 400$ px (resized to $256\times 256$) $400\times 400$ px (resized to $256\times 256$)
Number of Tiles 188,006 7,102
Label Availability Unlabeled Manual wall annotations
Data Splitting Not applicable 6-fold leave-one-area-out cross-validation
Figure 3: Spatial partitioning of the BudjBimArea dataset. Orange lines indicate annotated European historic dry-stone walls near Tae Rak (Lake Condah), Victoria, Australia. The number of data samples per area is as follows: Area 1 - 186, Area 2 - 1128, Area 3 - 1138, Area 4 - 1233, Area 5 - 1604, and Area 6 - 1813.
4.2 Results
Evaluation Metrics
For supervised evaluation, we assess model performance on dry-stone wall mapping using standard semantic segmentation metrics: Precision, Recall, F1 Score, and mean Intersection over Union (mIoU). These metrics are computed at the pixel level and averaged across individual image tiles to ensure a consistent and fine-grained evaluation.
Quantitative Result
Table 2: Stone wall mapping performances of ConvNet and ViT architectures. Evaluation metrics (i.e., Precision, Recall, F1 Score, and mIoU) are averaged over individual image tiles from the BudjBimArea dataset. The best performance for each backbone is highlighted in bold.
Method Backbone Param. Pre-train Dataset Pre. Rec. F1 mIoU
Random Init. ResNet-50-Siam. 46M - 74.7 77.0 75.8 61.1
Supervised ResNet-50-Siam. 46M ImageNet-1K 76.9 81.4 79.1 65.4
DINO (Caron et al., 2020) ResNet-50-Siam. 46M ImageNet-1K 75.7 74.3 75.0 60.0
DINO-MC (Wanyan et al., 2024) ResNet-50-Siam. 46M SeCo-100K 78.0 80.1 79.0 65.3
DINO-CV (ours) ResNet-50-Siam. 46M BudjBimLand. 79.4 83.0 81.2 68.3
Random Init. WRN-50-2-Siam. 138M - 75.6 77.7 76.6 62.1
Supervised WRN-50-2-Siam. 138M ImageNet-1K 76.4 82.5 79.3 65.7
DINO-MC (Wanyan et al., 2024) WRN-50-2-Siam. 138M SeCo-100K 79.0 79.2 79.1 65.4
DINO-CV (ours) WRN-50-2-Siam. 138M BudjBimLand. 79.8 83.0 81.4 68.6
Random Init. ViT-S/16-Siam. 42M - 77.1 73.0 75.0 60.0
Supervised ViT-S/16-Siam. 42M ImageNet-1K 74.2 82.4 78.1 64.0
DINO (Caron et al., 2020) ViT-S/16-Siam. 42M ImageNet-1K 76.8 81.1 78.9 65.2
DINO-MC (Wanyan et al., 2024) ViT-S/8-Siam. 42M SeCo-100K 77.4 80.4 78.9 65.1
DINO-CV (ours) ViT-S/16-Siam. 42M BudjBimLand. 77.4 83.0 80.1 66.8
Table 2 reports the supervised stone-wall mapping performance of three backbone architectures: ResNet-50, Wide-ResNet-50-2, and ViT-Small. The models are evaluated under different pre-training strategies: random initialization, supervised pre-training on ImageNet-1K, self-supervised pre-training on ImageNet-1K via DINO, SeCo-100K pre-training using DINO-MC, and our proposed DINO-CV on the BudjBimLandscape dataset. The results demonstrate that DINO-CV uniformly outperforms all other pre-training baselines across all model architectures. Specifically, models pre-trained with DINO-CV achieve the highest F1 scores (ResNet50: 81.2%, Wide ResNet50: 81.4% and ViT-small: 80.1%) and mIoU (ResNet50: 68.3%, Wide ResNet50: 68.6% and ViT-small: 66.8%), indicating superior ability to detect and localize dry-stone walls in complex terrain. These performance gains are evident over both supervised and self-supervised baselines, highlighting the benefit of domain-specific, cross-view self-supervised pre-training enabled by DINO-CV. Importantly, the improvements hold across both ConvNet and ViT backbones, suggesting the generalizability and robustness of the proposed pre-training strategy.
((a))
((b))
((c))
Figure 4: Test results (mIoU%) of each vision backbone on BudjBimArea dataset (area 1-6).
Figure 4 provides a detailed comparison of mIoU scores across six areas of the BudjBimArea dataset for each model backbone and pre-training strategy. Overall, DINO-CV achieves the highest or comparable performance relative to all baseline methods. For both ResNet-50 and Wide ResNet-50-2, DINO-CV shows notable improvements in regions with challenging visual characteristics, such as dense vegetation or ambiguous terrain boundaries (e.g., Area 3 and Area 6). ViT-S/16 also benefits significantly from DINO-CV, particularly in areas where other pre-training methods struggle. For example, in Area 2 and Area 4, DINO-CV with ViT-small backbone improves mIoU by 0.07–0.1 compared to random initialization. These results further support the robustness and generalizability of DINO-CV across diverse spatial regions of dataset and architectural designs.
Qualitative Result
Figure 5: Qualitative results of predictions using ResNet-50, Wide ResNet-50-2, and ViT-S/16 backbones under three pre-training strategies: random initialization, supervised training on ImageNet-1k, and DINO-CV (columns 1–3). Column 4 shows the ground truth (GT) segmentation masks, and column 5 provides the corresponding aerial imagery of the landscape.
Figs. 5 present qualitative results of dry-stone wall mapping using the ResNet-50, Wide ResNet-50-2, and ViT-S/16 backbones under three pre-training strategies: random initialization, supervised pre-training, and DINO-CV. Each figure illustrates the predicted segmentation masks for three representative areas from the BudjBimArea dataset.
The results demonstrate that models pre-trained with DINO-CV produce more accurate and spatially coherent segmentation masks compared to those initialized randomly or pre-trained via supervised learning. In Area 3 (Fig. 5–ResNet50), the ResNet-50 model with DINO-CV accurately captures intricate wall structures and subtle boundaries, whereas the randomly initialized baseline fails to detect substantial portions of the wall segments. Similarly, in Area 4 (Fig. 5–WideResNet50), the Wide ResNet-50-2 model with DINO-CV more effectively delineates complex wall boundaries in vegetated terrain, outperforming both the randomly initialized and supervised baselines, which misclassify several stone wall fragments. ViT-S/16 also shows notable qualitative improvements in Area 1 (Fig. 5–ViT-S/16), where DINO-CV enhances the integrity of parallel wall segmentation, compared to the fragmented predictions of the other baselines. These qualitative results align with the quantitative performance trends and highlight the effectiveness of DINO-CV in improving the mapping capacity of both ConvNet and ViT backbones across diverse and challenging landscape settings.
Generalization to Unseen Region
Figure 6: Generalization performance of the DINO-CV model with a ResNet-50-Siamese backbone on an unseen region within the Budj Bim Cultural Landscape. The upper panel shows the aerial imagery, with the blue rectangle marking the training area and the green rectangle denoting the unseen test area. The lower panel zooms into the unseen region, where the model accurately segments stone wall structures (annonated in white lines) beneath dense vegetation. These results highlight DINO-CV’s ability to learn transferable geomorphic representations from multi-view DEM derivatives (MHS and VAT), enabling large-scale mapping beyond the spatial extent of the training data.
To evaluate the spatial generalization of DINO-CV, the fine-tuned ResNet-50-Siamese model was tested on a region (green rectangle) located 10 km from the training area (blue rectangle), as shown in Figs. 6. Despite the absence of labeled samples, the model successfully delineates (in white) linear stone wall structures that are largely obscured by dense vegetation. This demonstrates robust transferability of learned representations across distinct spatial and ecological conditions. The cross-view self-supervision on DEM derivatives (MHS and VAT as inputs) enables the model to capture geomorphic patterns characteristic of archaeological stone walls. These results indicate that DINO-CV provides a scalable framework for large-scale, automated mapping of stone wall networks in complex and vegetated landscapes, reducing reliance on extensive labeled data, and supporting reliable mapping in areas that are inaccessible or hazardous for field survey.
5 Discussion
Table 3: Ablation study on the DINO-CV. We analyze the effectiveness of two key design components: (1) the effect of cross-view self-supervised pre-training of DINO-CV, and (2) the use of the Pseudo Siamese Architecture for supervised fine-tuning. Each row reports the F1 score and mean Intersection over Union (mIoU) for a specific configuration, evaluated on the BudjBimArea dataset. The columns $\Delta_{F1}$ and $\Delta_{mIoU}$ denote the performance difference relative to the DINO-CV baseline (row 1, 5, and 9 for each backbone).
Method Architecture Input Cross-view F1 $\Delta_{F1}$ mIoU $\Delta_{mIoU}$
1 DINO-CV ResNet-50-Siam. MHS+VAT 81.2 - 68.3 -
2 ResNet-50-Siam. MHS+VAT $\times$ 80.9 -0.3 67.9 -0.4
3 ResNet-50 MHS 80.8 -0.4 67.8 -0.5
4 ResNet-50 VAT 80.9 -0.3 68.0 -0.3
5 DINO-CV WRN-50-2-Siam. MHS+VAT 81.4 - 68.6 -
6 WRN-50-2-Siam. MHS+VAT $\times$ 81.0 -0.4 68.0 -0.6
7 WRN-50-2 MHS 80.8 -0.6 67.7 -0.9
8 WRN-50-2 VAT 81.0 -0.4 68.1 -0.5
9 DINO-CV ViT-S/16-Siam. MHS+VAT 80.1 - 66.8 -
10 ViT-S/16-Siam. MHS+VAT $\times$ 79.6 -0.5 66.2 -0.6
11 ViT-S/16 MHS 79.2 -0.9 65.6 -1.2
12 ViT-S/16 VAT 79.4 -0.7 65.8 -1.0
Table 4: Model mIoU (%) Performance on BudjBimArea training subsets.
Method Backbone 50% 30% 10%
Random Init. ResNet-50-Siam. 51.6 38.3 21.4
Supervised ResNet-50-Siam. 64.9 62.9 49.3
DINO-CV ResNet-50-Siam. 67.7 67.2 63.6
Random Init. WRN-50-2-Siam. 53.6 38.4 22.5
Supervised WRN-50-2-Siam. 65.3 63.6 49.7
DINO-CV WRN-50-2-Siam. 68.0 66.7 63.8
Random Init. ViT-S/16-Siam. 52.9 42.8 14.3
Supervised ViT-S/16-Siam. 63.2 61.6 56.4
DINO-CV ViT-S/16-Siam. 65.1 63.2 58.9
((a))
((b))
Figure 7: Self-attention maps from the pre-trained DINO-CV ViT-S/16 model without supervised fine-tuning.
((a))
((b))
Figure 8: Feature similarity maps of ViT and ResNet backbones on MHS and VAT inputs. MHS and VAT samples are randomly selected from the BudjBimArea dataset. The diagonal of each matrix indicates the cosine similarity between MHS and VAT features from the same location, reflecting the degree of cross-view alignment.
Ablation Study on DINO-CV
In Table 3, we conduct an ablation study to evaluate the effectiveness of two key design choices of our mapping approach: (1) the use of DINO-CV for pre-training, and (2) the Pseudo Siamese architecture for fusing representations from multiple DEM derivatives (MHS and VAT) during supervised fine-tuning.
Across all backbones (i.e., ResNet-50, WRN-50-2, and ViT-S/16), removing the cross-view component (rows 2, 6, 10) leads to consistent performance drops in both F1 and mIoU, confirming the effectiveness of cross-view training in learning view-invariant representations. This supports our core hypothesis that encouraging alignment across different DEM modalities enhances generalization and robustness. Furthermore, removing the Pseudo Siamese structure and instead using a single view (either MHS or VAT; rows 3–4, 7–8, 11–12) also results in lower performance compared to the dual-view configuration. The improvements from data view fusion demonstrate the complementary nature of MHS and VAT features, validating our architectural design for dual-stream representation learning. Together, these results show that both the cross-view pre-training and the dual-stream Pseudo Siamese architecture contribute positively and independently to model performance, with the full DINO-CV configuration achieving the best results across all backbones.
Learning View-invariant Features
To gain insight into the spatial attention patterns of models trained with DINO-CV, we visualize the self-attention maps produced by the ViT-S/16 backbone in Fig. 7, without any supervised fine-tuning. These maps highlight the regions of the input DEM derivatives (i.e., MHS and VAT) that the model attends to during inference. The visualizations reveal that DINO-CV enables the model to consistently focus on linear features resembling dry-stone walls. In particular, attention heads in intermediate ViT layers (i.e., attention layers 4 and 5) emphasize wall-like structures, while other attention layers (i.e., attention layer 1, 3 and 6) capture broader contextual information, such as the surrounding terrain and slope.
To further evaluate the alignment of feature representations across data views, we analyze the cosine similarity between the latent features extracted from MHS and VAT inputs using the teacher networks of DINO-CV. This analysis is conducted on both ViT and ResNet backbones, without any supervised fine-tuning. As shown in Fig. 8, DINO-CV exhibits substantially higher cosine similarity along the diagonal of the similarity matrix of data views compared to the original DINO (without cross-view learning), indicating improved feature alignment and enhanced view-invariance. These results confirm that DINO-CV effectively reduces inter-view feature discrepancies—particularly in mid-level spatial regions—and support the hypothesis that cross-view pre-training enables robust and consistent representations across diverse DEM derivatives.
Fine-tuning with Fewer Labels
Table 4 presents the mIoU performance of various models trained on different proportions (50%, 30%, and 10%) of annotated data from the BudjBimArea dataset. This scenario reflects real-world constraints, where annotation is costly and requires domain expertise, particularly in archaeological and heritage landscapes.
Across all backbone architectures, models pre-trained with DINO-CV consistently outperform both randomly initialized and ImageNet-supervised baselines, with the largest gains observed under low-label regimes. Using only 10% of annotated data, the DINO-CV pre-trained ResNet-50-Siamese model achieves 63.6% mIoU, compared to 49.3% for the supervised baseline and 21.4% for random initialization. A similar pattern is observed for the WRN-50-2-Siamese model, which achieves 63.8% mIoU at 10% supervision, outperforming the supervised and randomly initialized baselines by 14.1% and 41.3%, respectively.
The largest relative improvement is observed with the ViT-S/16-Siamese model. Under 10% supervision, the randomly initialized ViT collapses to 14.3% mIoU, while DINO-CV enables it to retain 58.9% mIoU, a gain of over 44% mIoU. Even compared to the supervised ViT (56.4%), DINO-CV improves performance across all levels of label availability.
These results highlight the value of cross-view self-supervised learning in enabling data-efficient feature learning for object mapping in archaeological landscapes. By leveraging unlabeled DEM derivatives, DINO-CV significantly reduces the reliance on large annotated datasets while maintaining high segmentation performance across both ConvNet and ViT backbones. This makes our approach especially suitable for large-scale, label-scarce tasks such as dry-stone wall mapping in cultural heritage environments.
6 Conclusions
In this study, we introduced DINO-CV, a cross-view self-supervised learning framework for mapping dry-stone walls in archaeologically significant and visually complex landscapes. Leveraging multi-view DEM derivatives of archaeological landscape, MHS and VAT, DINO-CV enables models to learn view-invariant and semantically rich features in self-supervised pre-training. Our method significantly improves stone wall mapping performance over both supervised and existing self-supervised baselines across ConvNet and ViT backbones. Under limited annotation, DINO-CV demonstrates strong label efficiency and generalization, making it highly suitable for large-scale cultural heritage mapping tasks.
Limitations
An inevitable limitation lies in the accuracy of wall annotations in the BudjBimArea dataset. Given the challenges in visually interpreting dense vegetation, shadowing, and subtle terrain features, some wall labels may suffer from spatial uncertainty or omission. Additionally, while the dataset focuses on a geographically and culturally rich region, its domain specificity may limit the direct transferability of results to other archaeological landscapes with different geomorphological characteristics or wall typologies.
Future Work
Future work could explore improving annotation quality of BudjBimArea dataset through semi-automated labeling strategies and integrating expert feedback loops. Additionally, extending the DINO-CV framework to support multi-modal data fusion, including hyperspectral imagery and spatial-temporal data, could potentially further enhance model’s mapping capability and robustness. Finally, applying DINO-CV to other cultural heritage sites worldwide for stone wall mapping could help assess its generalizability and scalability.
CRediT authorship contribution statement
Zexian Huang: Conceptualization, Writing – original draft, Writing – review & editing, Software, Methodology, Investigation. Mashnoon Islam: Conceptualization, Data curation. Brian Armstrong: Writing – review & editing. Kourosh Khoshelham: Writing – review & editing, Supervision. Martin Tomko: Writing – review & editing, Supervision.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
The source code and data support the findings of this study are available in following link: https://github.com/MLinArcheaomatics/BudjBimStoneWall-DINO-CV.
Acknowledgements
We thank Gunditj Mirring Traditional Owners Corporation (GMTOC) for providing the research data and overseeing its usage in this study, their invaluable cooperation, support, and meticulous efforts have significantly contributed to the quality of our research.